Draw the circle whose equation is given by 3x 2 + 3y 2 + 6x − 9y = 2.
1
Expert's answer
2020-10-14T11:01:42-0400

"3x ^2 + 3y ^2 + 6x − 9y = 2,\\3x ^2 + 3y^ 2 + 6x − 9y - 2=0,\\ x ^2 + y^ 2 + 2x − 3y - 2/3=0"


From here we can see that the center is "(-1,3\/2)" and radius is "\sqrt{1^2+1.5^2+2/3}=1.97"